# DyRoNet: Dynamic Routing and Low-Rank Adapters for Autonomous Driving Streaming Perception

### v3

@@ 4 Experiments
[…]
@@
Adaptive Streamer Ghosh et al. (2021) - 21.3 37.3 21.1 4.4 18.7 47.1
Streamer (S=600) Li et al. (2020) - 20.4 35.6 20.8 3.6 18.0 47.2
Streamer (S=900) Li et al. (2020) - 18.2 35.3 16.8 4.7 14.4 34.6
Streamer+AdaScale Ghosh et al. (2021) - 13.8 23.4 14.2 0.2 9.0 39.9
@@
DAMO-StreamNetNet-L He et al. (2023) 26.6 37.8 59.1 38.6 16.1 39.0 64.6
LongShortNet-L Li et al. (2023) 20.1 37.1 57.8 37.7 15.2 37.3 63.8
StreamYOLO-L Yang et al. (2022) 18.2 36.1 57.6 35.6 13.8 37.1 63.3
DAMO-StreamNetNet-M He et al. (2023) 24.3 35.7 56.7 35.9 14.5 36.3 63.3
LongShortNet-M Li et al. (2023) 17.5 34.1 54.8 34.6 13.3 35.3 58.1
StreamYOLO-M Yang et al. (2022) 18.2 32.9 54.0 32.5 12.4 34.8 58.1
DAMO-StreamNetNet-S He et al. (2023) 21.3 31.8 52.3 31.0 11.4 32.9 58.7
LongShortNet-S Li et al. (2023) 14.6 29.8 50.4 29.5 11.0 30.6 52.8
StreamYOLO-S Yang et al. (2022) 14.2 28.8 50.3 27.6 9.7 30.7 53.1
DyRoNet
DyRoNet ($\text{DAMO}_{\text{M + L}}$) 37.61 37.8 (+2.1) 58.8 (+2.1) 38.8 (+2.9) 16.1 (+1.6) 39.0 (+2.7) 64.0 (+0.7)
DyRoNet ($\text{LSN}_{\text{M + L}}$) 29.05 36.9 (+2.8) 58.2 (+3.4) 37.4 (+2.8) 14.9 (+1.6) 37.5 (+2.2) 63.3 (+5.2)
DyRoNet ($\text{sYOLO}_{\text{M + L}}$) 23.51 35.0 (+2.1) 55.7 (+1.7) 35.5 (+3.0) 13.7 (+1.3) 36.2 (+1.4) 61.1 (+3.0)
Table 1: The comparison of DyRoNet and SOTA. In this table, the optimal values are highlighted in green font and the online evaluation latency reaches the real-time is shown in red font.
@@ 4.1 Dataset and Metric
Dataset. For the evaluation of our methods, we utilized the comprehensive Argoverse-HD dataset Li et al. (2020), specifically designed for streaming perception in autonomous driving scenarios. This dataset comprises high-resolution RGB images captured from urban city street drives, offering a realistic representation of diverse driving conditions. The dataset is structured into two main segments: a training set consisting of 65 video clips and a test set comprising 24 video clips. Each video segment in the dataset, on average, spans over 600 frames, contributing to a training set with approximately 39k frames and a validation set containing around 15k frames. Notably, the Argoverse-HD dataset provides high-frame-rate (30fps) 2D object detection annotations, ensuring accuracy and reliability without relying on interpolated data.
Evaluation Metric. We adopt the streaming Average Precision (sAP) as the primary metric for performance evaluation. The sAP metric, widely recognized for its effectiveness in streaming perception tasks Li et al. (2020), offers a comprehensive assessment by calculating the mean Average Precision (mAP) across various Intersection over Union (IoU) thresholds, ranging from 0.5 to 0.95. This metric allows us to evaluate detection performance across different object sizes, including large, medium, and small objects, providing a robust measure of our model’s capability in real-world streaming perception scenarios.
@@ 4.2 Implementation Details
We tested three state-of-the-art streaming perception models: StreamYOLOYang et al. (2022), LongShortNetLi et al. (2023), and DAMO-StreamNetHe et al. (2023). These models, integral to the DyRoNet architecture, come with pre-trained parameters across three distinct scales: small (S), medium (M), and large (L), catering to a variety of processing requirements. In constructing the model bank $\mathcal{P}$ for DyRoNet, we strategically selected different model configurations to evaluate performance across diverse scenarios. For instance, the notation DyRoNet ($\text{DAMO}_{\text{S + M}}$) represents a configuration where DyRoNet employs the small (S) and medium (M) scales of DAMO-StreamNet as its two branches. [^3] Similar notations are used for other model combinations, allowing for a systematic exploration of the framework’s adaptability and performance under varying computational constraints. All experiments were conducted on a high-performance computing platform equipped with Nvidia 3090Ti GPUs (x4), ensuring robust and reliable computational power to handle the intensive processing demands of the streaming perception models. This setup provided a consistent and controlled environment for evaluating the efficacy of DyRoNet across different model configurations, contributing to the thoroughness and validity of our results. For more implementation details, please refer to Appendix C.
@@ 4.3 Comparision with SOTA Methods
We compared our proposed approach with state-of-the-art methods to evaluate its performance. In this subsection, we directly copied the reported performance from their original papers as their results. The performance comparison was conducted on the Argoverse-HD dataset Li et al. (2020). An overview of the results reveals that our proposed DyRoNet with a model bank of DAMO-StreamNet series achieves 37.8% sAP in 39.60 ms latency, outperforming the current state-of-the-art methods in latency by a significant margin. For the StreamYOLO and LongShortNet model banks, our DyRoNet attains 36.9% and 37.1% sAP in 29.35 ms, and 30.48 ms latency respectively, surpassing the original model dramatically. This demonstrates the effectiveness of the systematic improvements in DyRoNet.
Model Bank Random LoRA + Router
$\text{StreamYOLO}_{\text{S + M}}$ 39.16 26.25
$\text{StreamYOLO}_{\text{S + L}}$ 24.04 29.35
$\text{StreamYOLO}_{\text{M + L}}$ 24.69 23.51
$\text{LongShortNet}_{\text{S + M}}$ 24.79 21.47
$\text{LongShortNet}_{\text{S + L}}$ 21.49 30.48
$\text{LongShortNet}_{\text{M + L}}$ 24.75 29.05
@@
Table 2: Comparison of inference time (ms) on single RTX 3090. The optimal inference time between random and after train are consistently highlighted in green font.
Model Bank Full LoRA
$\text{StreamYOLO}_{\text{S + M}}$ 32.9 33.7
$\text{StreamYOLO}_{\text{S + L}}$ 36.1 36.9
$\text{StreamYOLO}_{\text{M + L}}$ 36.2 35.0
$\text{LongShortNet}_{\text{S + M}}$ 29.0 30.5
$\text{LongShortNet}_{\text{S + L}}$ 36.2 37.1
$\text{LongShortNet}_{\text{M + L}}$ 36.3 36.9
[…]
Table 3: Comparion of LoRA finetune and Full finetune. Full means the full fine-tuning and LoRA means the LoRA fine-tuning. And the best values between Full and LoRA are shown in red font.
Model Bank $b_{0}$ $b_{1}$ $b_{2}$ Random sAP
$K=2$ same model $\text{DAMO}_{\text{S + M}}$ 31.8 35.5 - 33.5 35.5
$\text{DAMO}_{\text{S + L}}$ 31.8 37.8 - 34.5 37.8
$\text{DAMO}_{\text{M + L}}$ 35.5 37.8 - 36.5 37.8
$\text{LSN}_{\text{S + M}}$ 29.8 34.1 - 31.8 30.5
$\text{LSN}_{\text{S + L}}$ 29.8 37.1 - 33.4 37.1
$\text{LSN}_{\text{M + L}}$ 34.1 37.1 - 35.6 36.9
$\text{sYOLO}_{\text{S + M}}$ 29.5 33.7 - 31.5 33.7
$\text{sYOLO}_{\text{S + L}}$ 29.5 36.9 - 33.2 36.9
$\text{sYOLO}_{\text{M + L}}$ 33.7 36.9 - 35.4 35.0
$K=2$ different model $\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{S}}$ 31.8 29.8 - 30.7 30.5
$\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{M}}$ 31.8 34.1 - 32.6 34.1
$\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{L}}$ 31.8 37.1 - 34.3 31.8
$\text{DAMO}_{\text{M}}$ + $\text{LSN}_{\text{S}}$ 35.5 29.8 - 32.6 29.8
$\text{DAMO}_{\text{L}}$ + $\text{LSN}_{\text{S}}$ 37.8 29.8 - 33.8 29.8
$K=3$ same model $\text{DAMO}_{\text{S + M + L}}$ 31.8 35.5 37.8 34.8 37.7
$\text{LSN}_{\text{S + M + L}}$ 29.8 34.1 37.1 33.5 36.1
$\text{sYOLO}_{\text{S + M + L}}$ 29.5 33.7 36.9 33.4 36.6
Table 4: Ablation of model bank setting. $K$ means the number of the model in bank $\mathcal{P}$.
@@ 4.4 Inference Time
We conducted detailed experiments analyzing the trade-offs between DyRoNet’s inference time and performance under different model bank selection strategies. Table 2 systematically presents the findings, with optimal times in green. This highlights DyRoNet’s superior performance—maintaining competitive inference speed alongside accuracy gains versus the random approach. Specifically, DyRoNet achieves efficient speeds while preserving or enhancing performance. This balance enables meeting real-time needs without compromising perception quality, critical for autonomous driving where both factors are paramount. By validating effectiveness in inference time reductions and accuracy improvements, the results show the practicality and efficiency of DyRoNet’s dynamic model selection.
@@ 4.5 Ablation Study
Router Network. To validate the effectiveness of the Router Network based on frame difference, we conducted comparative experiments using frame difference $\Delta I_{t}$, the current frame $I_{t}$, and the concatenation of the current frame with the previous historical frame $[I_{t}+I_{t-1}]$ as input modality of the Router Network. The experimental results are presented in Tab. 5. To control variables, in these experiments, we froze the model bank during training and only trained the Router Network. And only three different choices of StreamYOLO is involved in model bank. It can be obverse that using frame difference as input exhibits better performance than other two types of input modalities (35.0 of $\text{StreamYOLO}_{\text{S + L}}$ and 34.6 of $\text{StreamYOLO}_{\text{M + L}}$). This indicates that utilizing frame differences offers significant advantages in comprehending and characterizing environmental speed. Conversely, it also underscores that employing single frames as input or using multiple frames as input renders the lightweight model bank selection model ineffective.
Model Bank Input Modality LoRA
$\text{StreamYOLO}_{\text{S + M}}$ $I_{t}$ 33.7
$[I_{t}+I_{t-1}]$ 33.7
$\Delta I_{t}$ 32.6
$\text{StreamYOLO}_{\text{S + L}}$ $I_{t}$ 34.1
$[I_{t}+I_{t-1}]$ 30.2
$\Delta I_{t}$ 35.0
$\text{StreamYOLO}_{\text{M + L}}$ $I_{t}$ 33.7
$[I_{t}+I_{t-1}]$ 33.7
$\Delta I_{t}$ 34.6
@@
Table 5: Ablation of router network input. The optimal results are marked in red font under the same model bank setting.
@@
Branch Selection. Our research on streaming perception models has shown that configuring these models across varying scales can optimize their performance. We found that combining large and small models strikes an optimal balance, resulting in significant speed improvements. This conclusion is supported by the empirical evidence presented in Tab. 4, which clearly shows that the large-small model pairing outperforms both the large-small and large-medium combinations. Our findings highlight the importance of strategic model scaling in streaming perception and provide a framework for future model optimization in similar domains.
@@
Fine-tuning Scheme. In our evaluation, we contrasted the performance of direct fine-tuning with the Low-Rank Adapter (LoRA) fine-tuning strategy Zhu et al. (2023) for streaming perception models. Results are listed in Tab. 3. The results clearly demonstrated that LoRA fine-tuning surpasses direct fine-tuning, with the DAMO-Streamnet-based model bank configuration realizing an absolute gain of over 1.6%. This substantiates LoRA’s fine-tuning proficiency in circumventing the pitfalls of forgetting and data distribution bias inherent to direct fine-tuning. This experimental result demonstrates that LoRA fine-tuning can effectively mitigate the overfitting problem that may arise during model bank fine-tuning, leading to a stable and overall performance improvement.
@@
LoRA Rank. To assess the impact of different LoRA ranks in DyRoNet, we conducted experiments with $\mathrm{rank}=32,16,8$ respectively. All these experiments were set to train for 5 epochs, and the training alternated between Router Network training and model bank fine-tuning. The results are presented in Tab. 6. It can be observed that the performance is better with $rank=32$ compared to $rank=8$ and $rank=16$, and only occupy 10% of the total model parameters. Therefore, based on these experiments, $rank=32$ was selected as the default setting for our experiments. Although a smaller LoRA rank occupies fewer parameters, it leads to a rapid performance decay. The experimental results clearly demonstrate that with LoRA fine-tuning, it is possible to achieve superior performance than a single model while utilizing a smaller parameter footprint.
@@
$\text{DAMO}_{\text{S + L}}$ 16 31.8 37.8 35.9 7.73
$\text{DAMO}_{\text{S + L}}$ 8 31.8 37.8 35.9 4.02
@@
$\text{LSN}_{\text{S + L}}$ 16 29.8 37.1 30.6 5.48
$\text{LSN}_{\text{S + L}}$ 8 29.8 37.1 30.6 5.48
@@
$\text{sYOLO}_{\text{S + L}}$ 16 29.5 36.9 35.0 5.38
$\text{sYOLO}_{\text{S + L}}$ 8 29.5 36.9 35.0 2.7
Table 6: Ablation of LoRA rank: In the Param. column, we solely compare the proportion of parameters occupied by LoRA to the entire model. The best performance under the same model bank setting are highlighted in red font.

### v4

@@ 4 Experiments
Model Random MoE DyRoNet
Bank latency sAP latency sAP latency sAP sAP50 sAP75 sAPs sAPm sAPl
$\text{sYOLO}_{\text{S + M}}$ 39.16 31.5 66.16 29.5 26.25 (-12.91) 33.7 (+2.2) 53.9 34.1 13.0 35.1 59.3
$\text{sYOLO}_{\text{S + L}}$ 24.04 33.2 70.19 29.5 29.35 (+5.31) 36.9 (+3.7) 58.2 37.5 14.8 37.4 64.2
$\text{sYOLO}_{\text{M + L}}$ 24.69 35.4 83.65 33.7 23.51 (-1.18) 35.0 (-0.4) 55.7 35.5 13.7 36.2 61.1
$\text{LSN}_{\text{S + M}}$ 24.79 31.8 128.74 29.8 21.47 (-3.32) 30.5 (-1.3) 51.2 30.2 11.3 31.1 56.1
$\text{LSN}_{\text{S + L}}$ 21.49 33.4 121.62 29.8 30.48 (+8.99) 37.1 (+3.7) 58.3 37.6 15.1 37.6 63.7
$\text{LSN}_{\text{M + L}}$ 24.75 35.6 136.66 34.1 29.05 (+4.30) 36.9 (+1.3) 58.2 37.4 14.9 37.5 63.3
$\text{DAMO}_{\text{S + M}}$ 36.61 33.5 188.42 31.8 33.22 (-3.39) 35.5 (+2.0) 56.9 36.2 14.4 36.8 63.2
$\text{DAMO}_{\text{S + L}}$ 35.12 34.5 188.57 31.8 39.60 (+4.48) 37.8 (+3.3) 59.1 38.7 16.1 39.0 64.2
$\text{DAMO}_{\text{M + L}}$ 37.30 36.5 195.87 35.5 37.61 (+0.31) 37.8 (+1.3) 58.8 38.8 16.1 39.0 64.0
Table 1: Comparison of latency (ms) and the corresponding sAP performance on 1x RTX 3090. The values are highlighted in bold font if the DyRoNet perform better than the corresponding random and MoE case. Due to the overall poorer performance of MoE, we only lists the relative latency and sAP differences between DyRoNet and the random approach.
Methods Latency (ms) sAP $\uparrow$
@@
Adaptive Streamer [7] - 21.3
Streamer (S=600) [20] - 20.4
Streamer (S=900) [20] - 18.2
Streamer+AdaScale [7] - 13.8
@@
DAMO-StreamNet-L [10] 39.6 37.8
LongShortNet-L [19] 29.9 37.1
StreamYOLO-L [35] 29.3 36.1
DAMO-StreamNet-M [10] 33.5 35.7
LongShortNet-M [19] 25.1 34.1
StreamYOLO-M [35] 24.8 32.9
DAMO-StreamNet-S [10] 30.1 31.8
LongShortNet-S [19] 20.3 29.8
StreamYOLO-S [35] 21.3 28.8
DyRoNet ($\text{DAMO}_{\text{M + L}}$) 37.61 (-1.99) 37.8 (same)
DyRoNet ($\text{LSN}_{\text{M + L}}$) 29.05 (-0.85) 36.9 (-0.2)
DyRoNet ($\text{sYOLO}_{\text{M + L}}$) 23.51 (-5.79) 35.0 (-1.1)
Table 2: The comparison of DyRoNet and SOTA. The optimal values over its larger model are highlighted in bold font and the optimal values of online evaluation latency are shown in underline font. The latency of DyRoNet is evaluated on 1x RTX 3090 and compared with the latency of corresponding smaller model.
@@ 4.1 Dataset and Metric
Dataset. Argoverse-HD dataset [20] is utilized for our experiments, specifically designed for streaming perception in autonomous driving scenarios. It comprises high-resolution RGB images captured from urban city street, offering a realistic representation of diverse driving conditions. The dataset is structured into two main segments: a training set consisting of 65 video clips and a test set comprising 24 video clips. Each video clip in the dataset spans over 600 frames in average, contributing to a training set with approximately 39k frames and a test set containing around 15k frames. Notably, Argoverse-HD provides high-frame-rate (30fps) 2D object detection annotations, ensuring accuracy and reliability without relying on interpolated data.
Evaluation Metric. Streaming Average Precision (sAP) are adopted as the primary metric for performance evaluation, which is widely recognized for its effectiveness in streaming perception tasks [20]. It offers a comprehensive assessment by calculating the mean Average Precision (mAP) across various Intersection over Union (IoU) thresholds, ranging from 0.5 to 0.95. This metric allows us to evaluate detection performance across different object sizes, including large, medium, and small objects, providing a robust measurement in real-world streaming perception scenarios.
@@ 4.2 Implementation Details
We tested three state-of-the-art streaming perception models: StreamYOLO[35], LongShortNet[19], and DAMO-StreamNet[10]. These models, integral to the DyRoNet architecture, come with pre-trained parameters across three distinct scales: small (S), medium (M), and large (L), catering to a variety of processing requirements. In constructing the model bank $\mathcal{P}$ for DyRoNet, we strategically selected different model configurations to evaluate performance across diverse scenarios. For instance, the notation DyRoNet ($\text{DAMO}_{\text{S + M}}$) represents a configuration where DyRoNet employs the small (S) and medium (M) scales of DAMO-StreamNet as its two branches. [^3] Similar notations are used for other model combinations, allowing for a systematic exploration of the framework’s adaptability and performance under varying computational constraints. All experiments were conducted on a high-performance computing platform equipped with Nvidia 3090Ti GPUs (x4), ensuring robust and reliable computational power to handle the intensive processing demands of the streaming perception models. This setup provided a consistent and controlled environment for evaluating the efficacy of DyRoNet across different model configurations, contributing to the thoroughness and validity of our results. For more implementation details, please refer to Appendix C.
@@ 4.3 Comparision with SOTA Methods
We compared DyRoNet with state-of-the-art methods to evaluate its performance. In this subsection, we directly copied the reported sAP performance from their original papers as their results, for fair comparison, we evaluated the latency of each real-time model on 1x RTX 3090. The performance comparison was conducted on the Argoverse-HD dataset [20]. An overview of the results reveals that our proposed DyRoNet with a model bank of DAMO-StreamNet series achieves 37.8% sAP in 37.61 ms latency, outperforming the current state-of-the-art methods in latency by a significant margin. This demonstrates the effectiveness of the systematic improvements in DyRoNet.
Model Bank Full (sAP) LoRA (sAP)
$\text{StreamYOLO}_{\text{S + M}}$ 32.9 33.7 (+0.8)
$\text{StreamYOLO}_{\text{S + L}}$ 36.1 36.9 (+0.8)
$\text{StreamYOLO}_{\text{M + L}}$ 36.2 35.0 (-1.2)
$\text{LongShortNet}_{\text{S + M}}$ 29.0 30.5 (+1.5)
$\text{LongShortNet}_{\text{S + L}}$ 36.2 37.1 (+0.9)
$\text{LongShortNet}_{\text{M + L}}$ 36.3 36.9 (+0.6)
@@
Table 3: Comparison of LoRA finetune and Full finetune. The optimal values between Full and LoRA are shown in bold font.
Model Bank $b_{0}$ $b_{1}$ $b_{2}$ sAP
$K=2$ same model $\text{DAMO}_{\text{S + M}}$ 31.8 35.5 - 35.5
$\text{DAMO}_{\text{S + L}}$ 31.8 37.8 - 37.8
$\text{DAMO}_{\text{M + L}}$ 35.5 37.8 - 37.8
$\text{LSN}_{\text{S + M}}$ 29.8 34.1 - 30.5
$\text{LSN}_{\text{S + L}}$ 29.8 37.1 - 37.1
$\text{LSN}_{\text{M + L}}$ 34.1 37.1 - 36.9
$\text{sYOLO}_{\text{S + M}}$ 29.5 33.7 - 33.7
$\text{sYOLO}_{\text{S + L}}$ 29.5 36.9 - 36.9
$\text{sYOLO}_{\text{M + L}}$ 33.7 36.9 - 35.0
$K=2$ different model $\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{S}}$ 31.8 29.8 - 30.5
$\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{M}}$ 31.8 34.1 - 34.1
$\text{DAMO}_{\text{S}}$ + $\text{LSN}_{\text{L}}$ 31.8 37.1 - 31.8
$\text{DAMO}_{\text{M}}$ + $\text{LSN}_{\text{S}}$ 35.5 29.8 - 29.8
$\text{DAMO}_{\text{L}}$ + $\text{LSN}_{\text{S}}$ 37.8 29.8 - 29.8
$K=3$ same model $\text{DAMO}_{\text{S + M + L}}$ 31.8 35.5 37.8 37.7
$\text{LSN}_{\text{S + M + L}}$ 29.8 34.1 37.1 36.1
$\text{sYOLO}_{\text{S + M + L}}$ 29.5 33.7 36.9 36.6
Table 4: The performance of DyRoNet under various model bank selection. $K$ means the number of the model in bank $\mathcal{P}$.
@@ 4.4 Inference Time
In this subsection, we conducted detailed experiments analyzing the trade-offs between DyRoNet’s inference time and performance under different model bank selection. It is notable that the latency of DyRoNet presented in Tab. 2 do not accurately reflect the real outperform. Due to the varying hardware platforms for measuring latency across methods, a fair comparison cannot be achieved. For instance, DAMO-StreamNet is tested on 1x V100. To address these differences, we conducted additional tests on a 1x RTX 3090, which highlight DyRoNet’s performance enhancements. Tab. 1 presents the latency comparison and highlights DyRoNet’s superior performance—maintaining competitive inference speed alongside accuracy gains versus the random and MoE approaches. Where the MoE predicts the weights of each branch via a gate module and then combines the output results accordingly. Specifically, DyRoNet achieves efficient speeds while preserving or enhancing performance. This balance enables meeting real-time needs without compromising perception quality, critical for autonomous driving where both factors are paramount. By validating effectiveness in inference time reductions and accuracy improvements, the results show the practicality and efficiency of DyRoNet’s dynamic model selection.
@@ 4.5 Ablation Study
Router Network. To validate the effectiveness of the Router Network based on frame difference, we conducted comparative experiments using frame difference $\Delta I_{t}$, the current frame $I_{t}$, and the concatenation of the consecutive frames $[I_{t}+I_{t-1}]$ as input modality of the Router Network.  For comparison, a naive method of input guidance also employed: By calculating $\mathbb{E}(\Delta I_{t})$, the larger branch is selected if $\mathbb{E}(\Delta I_{t})>0$, otherwise the smaller branch is chosen.  The results are presented in Tab. 6. For comparison, the Router Network only be trained in these experiments while the leftover parts be frozen.
It shows that using $\Delta I_{t}$ as input exhibits better performance than other methods (35.0 sAP of $\text{sYOLO}_{\text{S + L}}$ and 34.6 sAP of $\text{sYOLO}_{\text{M+L}}$). This indicates that utilizing $\Delta I_{t}$ offers significant advantages in comprehending and characterizing environmental speed. Conversely, it also underscores that employing single frames as input or using multiple frames as input renders the lightweight model bank selection model ineffective. Furthermore, the proportion of sample splits across branches can also illustrate the discriminative power with respect to environmental factors. For instance, the $\mathbb{E}(\Delta I_{t})$ criterion resulting in a evenly spliting distribution (48.22% $\mathbb{E}(\Delta I_{t})>0$ over train set and 49.85% $\mathbb{E}(\Delta I_{t})>0$ over test set). Indicating the direct sample selection without router lacks estimation of environmental factors, thereby weakening its discriminative power.
In contrast, Tab.5 presents statistics indicating the router layer’s effectiveness in allocating samples to specific models and showcase its ability to strike a balance between latency and performance. This balance is crucial for streaming perception and underscores our contribution.
Model training time inference time
Combination Model 1 Model 2 Model 1 Model 2
SYOLO (M+L) 37.53% 62.47% 94.67% 5.33%
LSN (M+L) 30.86% 69.14% 19.87% 80.13%
DAMO (M+L) 84.61% 15.39% 0.02% 99.98%
Table 5: The statistics of model selection by DyRoNet under different model choices during both training and inference time.
Model Input Modality / Criterion
Bank $I_{t}$ $[I_{t}+I_{t-1}]$ $\mathbb{E}(\Delta I_{t})$ $\Delta I_{t}$ (DyRoNet)
$\text{sYOLO}_{\text{S + M}}$ 33.7 33.7 31.5 32.6
$\text{sYOLO}_{\text{S + L}}$ 34.1 30.2 32.9 35.0
$\text{sYOLO}_{\text{M + L}}$ 33.7 33.7 34.2 34.6
@@
Table 6: Ablation of router network input / criterion. The optimal results are marked in bold font under the same model bank setting.
@@
Branch Selection. Our research on streaming perception models has shown that configuring these models across varying scales can optimize their performance. We found that combining L and S models strikes an optimal balance, resulting in significant speed improvements. This conclusion is supported by the empirical evidence presented in Tab. 4, which clearly shows that the L+S model pairing outperforms both the L+S and L+M cases. Our findings highlight the importance of strategic model scaling in streaming perception and provide a framework for future model optimization in similar domains.
@@
Fine-tuning Scheme. We contrasted the performance of direct fine-tuning with the LoRA fine-tuning strategy [38] for streaming perception models.  Tab. 3 shows that LoRA fine-tuning surpasses direct fine-tuning, with the DAMO-Streamnet-based model bank configuration realizing an absolute gain of over 1.6%. This substantiates LoRA’s fine-tuning proficiency in circumventing the pitfalls of forgetting and data distribution bias inherent to direct fine-tuning. This result demonstrates that LoRA fine-tuning can effectively mitigate the overfitting while fine-tuning, leading to a stable performance improvement.
@@
LoRA Rank. To assess the impact of LoRA ranks in DyRoNet, we conducted experiments with rank $r=32,16,8$ respectively. All experiments were train for 5 epochs between Router Network training and model bank fine-tuning. The results are presented in Tab. 7. It can be observed that the performance is better with $r=32$ compared to $r=8$ and $r=16$, and only occupy 10% of the total model parameters. Therefore, based on these experiments, $r=32$ was selected as the default setting for our experiments. Although a smaller LoRA rank occupies fewer parameters, it leads to a rapid performance decay. The experimental results clearly demonstrate that with LoRA fine-tuning, it is possible to achieve superior performance than a single model while utilizing a smaller parameter footprint.
@@
$\text{DAMO}_{\text{S + L}}$ 8 31.8 37.8 35.9 4.02
$\text{DAMO}_{\text{S + L}}$ 16 31.8 37.8 35.9 7.73
@@
$\text{LSN}_{\text{S + L}}$ 8 29.8 37.1 30.6 5.48
$\text{LSN}_{\text{S + L}}$ 16 29.8 37.1 30.6 5.48
@@
$\text{sYOLO}_{\text{S + L}}$ 8 29.5 36.9 35.0 2.7
$\text{sYOLO}_{\text{S + L}}$ 16 29.5 36.9 35.0 5.38
@@
Table 7: Ablation of LoRA rank: In the Param. column, we solely compare the proportion of parameters occupied by LoRA to the entire model. The best performance under the same model bank setting are highlighted in bold font.
4.6 Extra Experiment on NuScenes-H dataset
To validate the DyRoNet on other dataset, we converted the 3D streaming perception dataset nuScenes-H[30] into 2D format. The experiment details are provided in the Appendix D. As shown in Tab. 8, DyRoNet consistently achieves better results than other branch fusion methods on nuScenes-H 2D dataset. It demonstrates DyRoNet’s advantages in branch fusion selection and its versatility.
Model Bank b1 b2 Random MoE DyRoNet
$\text{sYOLO}_{\text{S + M}}$ 6.9 9.2 8.1 6.9 8.9
$\text{sYOLO}_{\text{S + L}}$ 7.3 9.9 8.6 7.3 9.0
$\text{sYOLO}_{\text{M + L}}$ 8.9 9.6 9.2 8.9 9.3
$\text{sYOLO}_{\text{S}}+\text{LSN}_{\text{S}}$ 6.6 6.2 6.4 6.2 6.5
$\text{sYOLO}_{\text{M}}+\text{LSN}_{\text{M}}$ 9.1 9.3 9.2 9.1 9.3
$\text{sYOLO}_{\text{L}}+\text{LSN}_{\text{L}}$ 9.0 9.6 9.3 9.0 9.6
Table 8: The sAP results are shown in the table. Where b1 denotes the independent performance of the first branch and b2 denotes the second one. The branch fusion method Random and MoE are similar with Tab. 1. The best method is highlighted in bold font.